A flywheel of radius 10 inches is turning at the rate 900 rpm. How fast does a point on the rim travel in ft/sec? Answer by mananth(16949) 
 10 inches = 10/12 feet
= 0.83 feet
diameter = 1.66 feet
Circumference = pi*d
3.14 * 1.66 feet= 5.21 feet
rpm = 900 In 1 minute the point travels
5.21*900 =4691 feet
4691 feet in 60 sec
so 4691/60 = 78.2 ft/sec